OpenStudy (cutiecomittee123):

if the point p(-3/5,y) is a point on the unit circle in the 3rd quadrant what is y? I am pretty sure you need x^2+y^2=1^2

OpenStudy (cutiecomittee123):

https://www.wolframalpha.com/input/?i=-3%2F5%5E2%2Bx%5E2%3D1%5E2+solve+for+x+ I need a fraction answer and wolfram alpha is giving me a mixed number

OpenStudy (anonymous):

Since it's in third quadrant, the point would be (-x, -y), meaning the y-value is a negative. https://www.wolframalpha.com/input/?i=%28-3%2F5%29%5E2%2By%5E2%3D1%5E2+solve+for+y

OpenStudy (anonymous):

"I need a fraction answer and wolfram alpha is giving me a mixed number " it's because you put "-3/5^2+x^2=1^2 solve for x" , it should be "(-3/5)^2+y^2=1^2 solve for y." :D

OpenStudy (cutiecomittee123):

okay thanks